### v1

@@ Abstract
In this work, we systematically study the interactions of the $S$-wave $D^{(*)}\bar{B}^{(*)}$ systems within the framework of chiral effective field theory in heavy hadron formalism. We calculate the $D^{(*)}\bar{B}^{(*)}$ effective potentials up to next-to-leading order, explore the bound state formations, and investigate the $D^{(*)}\bar{B}^{(*)}$ scattering properties such as scattering rate, scattering length, and effective range. Our results show that all $I=1$ $D^{(*)}\bar{B}^{(*)}$ potentials are repulsive, preventing the formation of bound states, while the $I=0$ potentials are generally attractive. Specifically, we get two important observations: first, the shallow bound state is more likely to exist in the $D\bar{B}[I(J^{P})=0(0^{+})]$ system than in the $D\bar{B}^{*}[I(J^{P})=0(1^{+})]$ system; second, $D^{*}\bar{B}^{*}[I(J^{P})=0(0^{+})]$ and $D^{*}\bar{B}^{*}[I(J^{P})=0(1^{+})]$ systems possess relatively large binding energies and positive scattering lengths, which suggests strong bound state formations in these channels. So the attractions in the $D^{*}\bar{B}^{*}[I=0]$ systems are deeper than those in the $D\bar{B}^{(*)}[I=0]$ systems, thus we strongly recommend the future experiment to search for the $D^{*}\bar{B}^{*}[I=0]$ tetraquark systems. In addition, we also investigate the dependencies of the $D\bar{B}^{(*)}$ binding energies on the contact low-energy coupling constants (LECs).
@@ I Introduction
These experimental discoveries have spurred extensive theoretical investigations into the nature of open heavy-flavor multiquark states. Various approaches, including QCD sum rules, lattice QCD, and effective field theories, have been employed to explore their properties and structures (see Refs. Cui:2006mp ; Detmold:2007wk ; Carlucci:2007um ; Navarra:2007yw ; Yang:2009zzp ; Molina:2010tx ; Carames:2011zz ; Du:2012wp ; Hyodo:2012pm ; Ohkoda:2012hv ; Vijande:2013qr ; Ikeda:2013vwa ; Bicudo:2015kna ; Bicudo:2015vta ; Peters:2015tra ; Peters:2016isf ; Francis:2016hui ; Bicudo:2017szl ; Chen:2017rhl ; Azizi:2018mte ; Bicudo:2019mny ; Lu:2020qmp ; He:2020jna ; Cheng:2020nho ; Albuquerque:2020ugi ; Guo:2021mja ; Chen:2022asf for comprehensive reviews). These studies have deepened our understanding of the exotic hadron spectrum and highlighted the importance of investigating the interactions of dimeson systems, particularly those involving doubly heavy quarks.
@@
In this work, we investigate the interactions of $S$-wave $D\bar{B}$, $D\bar{B}^{*}$, and $D^{*}\bar{B}^{*}$ systems using chiral effective field theory (ChEFT) in the heavy hadron formalism. ChEFT, a low-energy effective theory of QCD, has proven to be a powerful tool for studying hadronic interactions, particularly in the context of nucleon-nucleon systems Machleidt:2011zz ; Hammer:2019poc . By incorporating heavy-quark and chiral symmetries, ChEFT provides a robust framework for analyzing the interactions of heavy mesons. Previous studies have applied ChEFT to systems such as $D^{(*)}D^{(*)}$ Xu:2017tsr ; Wang:2022jop , $\bar{B}^{(*)}\bar{B}^{(*)}$ Wang:2018atz , and hidden-charm systems Xu:2021vsi , as well as to the interactions of heavy pentaquark molecular states Meng:2019ilv ; Wang:2019ato ; Wang:2022ztm . These investigations have demonstrated the effectiveness of ChEFT in describing the low-energy dynamics of heavy hadrons.
This paper is organized as follows: In Sec. II, we introduce the chiral effective Lagrangians with $SU(2)$ flavor symmetry used in our calculations. In Sec. III, we present the scattering amplitudes and derive the effective meson-meson potentials. In Sec. IV, we analyze the numerical results, examining the behaviors of the effective potentials and searching for possible bound states by solving the Schrödinger equation. We also discuss the scattering $T$-matrix and extract physical quantities such as the scattering rate, scattering length, and effective range. Finally, in Sec. V, we summarize our findings and discuss their implications for future studies of doubly heavy tetraquark systems.
@@ II.2 Lagrangians at next-to-leading order
The low-energy constants appearing in the above equations are split into finite and infinite parts. The infinite parts are used to cancel the divergences of the loop contributions at NLO. The finite parts also contribute to the amplitudes, and they additionally introduce a large number of LECs.
@@
where $\mathcal{F}(\mathbf{q})$ is a regulator function. To avoid ultraviolet (UV) divergence in the integral, we use a Gaussian regulator function $\mathcal{F}(\mathbf{q})=\exp(-\mathbf{q}^{2n}/\Lambda^{2n})$ with $n=2$ Xu:2017tsr ; Wang:2018atz ; Xu:2021vsi . Based on the potentials in coordinate space, we can solve the Schrödinger equations to search for the possible bound states.
@@ III.1 $D\bar{B}$ systems
In the above amplitudes (19)-(29), the loop functions $J^{a/b}_{ij}(m,\omega)$, $J^{g/h}_{ij}(m,\omega_{1},\omega_{2})$, $J^{F}_{ij}(m_{1},m_{2},q)$, $J^{T/S}_{ij}(m_{1},m_{2},\omega,q)$, and $J^{B/R}_{ij}(m_{1},m_{2},\omega_{1},\omega_{2},q)$ are abbreviated as $J^{a/b}_{ij}$, $J^{g/h}_{ij}$, $J^{F}_{ij}$, $J^{T/S}_{ij}$, and $J^{B/R}_{ij}$, respectively. The variables $m,m_{1},m_{2}$ denote the pion mass, while the mass-splitting-dependent variables $\omega_{1}$ and $\omega_{2}$ for every amplitudes are collected in Tables 1-2. Here we define $\delta_{1}=M_{D^{*}}-M_{D}$ and $\delta_{2}=M_{\bar{B}^{*}}-M_{\bar{B}}$ as the $D^{*}$-$D$ and $\bar{B}^{*}$-$\bar{B}$ mass differences, respectively. The calculating procedures of these loop functions follow the Refs. Xu:2017tsr ; Xu:2021vsi .
Figure 1: LO contact, NLO contact and NLO TPE diagrams of the process $D\bar{B}\to D\bar{B}$. The thin solid, double thin solid, solid, thick solid, and dashed lines stand for the $\bar{B}$, $\bar{B}^{*}$, $D$, $D^{*}$ and a pion, respectively. TPE diagrams at NLO.
@@ III.2 $D\bar{B}^{*}$ systems
Figure 2: LO contact, NLO contact and NLO TPE diagrams of the process $D\bar{B}^{*}\to D\bar{B}^{*}$. The thin solid, double thin solid, solid, thick solid, and dashed lines stand for the $\bar{B}$, $\bar{B}^{*}$, $D$, $D^{*}$ and a pion, respectively.
@@
$\mathcal{M}_{a2.2}^{(2)}=$ $~{}4(d-3)(d-2)\frac{g_{2}^{2}}{f^{2}}AJ_{22}^{g}(\epsilon_{2}\cdot\epsilon^{*}_{4}),$ (33)
$\mathcal{M}_{a2.3}^{(2)}=$ $~{}4(d-1)\frac{g_{1}^{2}}{f^{2}}AJ_{22}^{g}(\epsilon_{2}\cdot\epsilon^{*}_{4}),$ (34)
@@
$\mathcal{M}_{a2.6}^{(2)}=$ $-4(d-3)(d-2)\frac{g_{1}g_{2}}{f^{2}}AJ_{22}^{h}(\epsilon_{2}\cdot\epsilon^{*}_{4}),$ (37)
$\mathcal{M}_{a2.7}^{(2)}=$ $-4(d-3)(d-2)\frac{g_{1}g_{2}}{f^{2}}AJ_{22}^{h}(\epsilon_{2}\cdot\epsilon^{*}_{4}),$ (38)
@@
$\mathcal{M}_{a2.10}^{(2)}=$ $~{}\frac{3}{2}(d-2)\frac{g_{1}^{2}}{2f^{2}}A\partial_{\omega}J_{22}^{b}(\epsilon_{2}\cdot\epsilon^{*}_{4}).$ (40)
@@
$\mathcal{M}_{c2.2}^{(2)}=$ $i4\frac{g_{2}^{2}}{f^{4}}\{[A_{1}(q_{0}J_{22}^{S}+J_{24}^{S}+q_{0}J_{32}^{S}+J_{33}^{S})-A_{5}(q_{0}J_{11}^{S}+2q_{0}J_{22}^{S}+J_{24}^{S}+q_{0}J_{32}^{S}+J_{33}^{S})](q\cdot\epsilon_{2})(q\cdot\epsilon^{*}_{4})$
@@
$\mathcal{M}_{c2.5}^{(2)}=$ $-iA_{1}4i\frac{g_{1}^{2}g_{2}^{2}}{f^{4}}\{[J_{21}^{B}-\vec{q}^{2}J_{22}^{B}+(d+3)(J_{31}^{B}+J_{42}^{B})-2\vec{q}^{2}J_{32}^{B}-\vec{q}^{2}J_{43}^{B}](q\cdot\epsilon_{2})(q\cdot\epsilon^{*}_{4})-[\vec{q}^{2}(J_{31}^{B}$
$+J_{42}^{B})-(1+d)J_{41}^{B}]\}(\epsilon_{2}\cdot\epsilon^{*}_{4}),$ (45)
$\mathcal{M}_{c2.6}^{(2)}=$ $-iA_{1}4i\frac{g_{1}^{2}g_{2}^{2}}{f^{4}}(d-3)\{[J_{21}^{B}-\vec{q}^{2}J_{22}^{B}+(d+3)J_{31}^{B}-2\vec{q}^{2}J_{32}^{B}+(d+3)J_{42}^{B}-\vec{q}^{2}J_{43}^{B}](q\cdot\epsilon_{2})(q\cdot\epsilon^{*}_{4})$
$+[-\vec{q}^{2}J_{21}^{B}+\vec{q}^{4}J_{22}^{B}-(2d+1)\vec{q}^{2}(J_{31}^{B}+J_{42}^{B})+2\vec{q}^{4}J_{32}^{B}+(d+1)(d-2)J_{41}^{B}+\vec{q}^{4}J_{43}^{B}]\}(\epsilon_{2}\cdot\epsilon^{*}_{4}),$ (46)
$\mathcal{M}_{c2.7}^{(2)}=$ $-iA_{1}4i\frac{g_{1}^{2}g_{2}^{2}}{f^{4}}\{[J_{21}^{R}-\vec{q}^{2}J_{22}^{R}+(d+3)(J_{31}^{R}+J_{42}^{R})-2\vec{q}^{2}J_{32}^{R}-\vec{q}^{2}J_{43}^{R}](q\cdot\epsilon_{2})(q\cdot\epsilon^{*}_{4})-[\vec{q}^{2}J_{31}^{R}$
$-(1+d)J_{41}^{R}+\vec{q}^{2}J_{42}^{R}]\}(\epsilon_{2}\cdot\epsilon^{*}_{4}),$ (47)
$\mathcal{M}_{c2.8}^{(2)}=$ $-i4i\frac{g_{1}^{2}g_{2}^{2}}{f^{4}}(d-3)A_{1}\{[J_{21}^{R}-\vec{q}^{2}J_{22}^{R}+(d+3)J_{31}^{R}-2\vec{q}^{2}J_{32}^{B}+3J_{42}^{B}-\vec{q}^{2}J_{43}^{B}](q\cdot\epsilon_{2})(q\cdot\epsilon^{*}_{4})+[-\vec{q}^{2}J_{21}^{R}$
$+\vec{q}^{4}J_{22}^{R}-(2d+1)\vec{q}^{2}(J_{31}^{R}+J_{42}^{R})+2\vec{q}^{4}J_{32}^{R}+(d+1)(d-2)J_{41}^{R}+\vec{q}^{4}J_{43}^{R}]\}(\epsilon_{2}\cdot\epsilon^{*}_{4}),$ (48)
@@
. $I=1$ $I=0$ $A$ $A$ $\omega_{1}$ $\omega_{2}$ $A_{a2.1}$ $\frac{1}{4}(3D_{a}-E_{a})$ $\frac{3}{4}(D_{a}+E_{a})$ $\delta_{2}$ $\delta_{2}$ $A_{a2.2}$ $\frac{1}{4}(3D_{a}-E_{a})$ $\frac{3}{4}(D_{a}+E_{a})$ 00 00 $A_{a2.3}$ 00 00 $-\delta_{1}$ $-\delta_{1}$ $A_{a2.4}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{-3}{4}(D_{b}-3E_{b})$ $\delta_{2}$ $-\delta_{1}$ $A_{a2.5}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{-3}{4}(D_{b}-3E_{b})$ $-\delta_{1}$ $\delta_{2}$ $A_{a2.6}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{-3}{4}(D_{b}-3E_{b})$ $\delta_{2}$ $-\delta_{1}$ $A_{a2.7}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{-3}{4}(D_{b}-3E_{b})$ $-\delta_{1}$ $\delta_{2}$ $A_{a2.(8+9)}$ $D_{a}+E_{a}$ $D_{a}-3E_{a}$ 00 $\delta_{2}$ $A_{a2.10}$ $D_{a}+E_{a}$ $D_{a}-3E_{a}$ $\delta_{1}$ 00
@@
       $A_{c2.2}$        $\frac{i}{8}$        $\frac{i}{8}$        00        $-\frac{3i}{8}$        $\frac{3i}{8}$        00        $\delta_{2}$        00
       $A_{c2.2}$        $\frac{i}{8}$        $\frac{i}{8}$        00        $-\frac{3i}{8}$        $\frac{3i}{8}$        00        $\delta_{2}$        00
       $A_{c2.3}$        $\frac{i}{8}$        $-\frac{i}{8}$        00        $-\frac{3i}{8}$        $\frac{3i}{8}$        00        00        00
@@
       $A_{c2.5}$        $\frac{1}{16}$        00        00        $\frac{9}{16}$        00        00        $-\delta_{1}$        $\delta_{2}$
       $A_{c2.6}$        $\frac{1}{16}$        00        00        $\frac{9}{16}$        00        00        $-\delta_{1}$        00
       $A_{c2.7}$        $\frac{5}{16}$        00        00        $-\frac{3}{16}$        00        00        $-\delta_{1}$        $\delta_{2}$
       $A_{c2.8}$        $\frac{5}{16}$        00        00        $\frac{-3}{16}$        00        00        $-\delta_{1}$        00

### v2

@@ Abstract
In this work, we systematically study the interactions of the $S$-wave $D^{(*)}\bar{B}^{(*)}$ systems within the framework of chiral effective field theory in heavy hadron formalism. We calculate the $D^{(*)}\bar{B}^{(*)}$ effective potentials up to next-to-leading order, explore the bound state formations, and investigate the $D^{(*)}\bar{B}^{(*)}$ scattering properties such as scattering rate, scattering length, and effective range. Our results show that all $I=1$ $D^{(*)}\bar{B}^{(*)}$ potentials are repulsive, preventing the formation of bound states, while the $I=0$ potentials are generally attractive. Specifically, we get two important observations: first, the shallow bound state is more likely to exist in the $D\bar{B}[I(J^{P})=0(0^{+})]$ system than in the $D\bar{B}^{*}[I(J^{P})=0(1^{+})]$ system; second, $D^{*}\bar{B}^{*}[I(J^{P})=0(0^{+})]$ and $D^{*}\bar{B}^{*}[I(J^{P})=0(1^{+})]$ systems possess relatively large binding energies and positive scattering lengths, which suggests strong bound state formations in these channels. So the attractions in the $D^{*}\bar{B}^{*}[I=0]$ systems are deeper than those in the $D\bar{B}^{(*)}[I=0]$ systems, thus we strongly recommend the future experiment to search for the $D^{*}\bar{B}^{*}[I=0]$ tetraquark systems. We also investigate the coupled-channel effects on the $J=0,1$ systems and conclude that the inclusion of the coupled channels introduces small but visible influences. In addition, we also investigate the dependencies of the $D\bar{B}^{(*)}$ binding energies on the contact low-energy coupling constants (LECs).
@@ I Introduction
These experimental discoveries have spurred extensive theoretical investigations into the nature of open heavy-flavor multiquark states. Various approaches, including QCD sum rules, lattice QCD, and effective field theories, have been employed to explore their properties and structures (see Refs. Cui:2006mp ; Detmold:2007wk ; Carlucci:2007um ; Navarra:2007yw ; Yang:2009zzp ; Molina:2010tx ; Carames:2011zz ; Du:2012wp ; Hyodo:2012pm ; Ohkoda:2012hv ; Vijande:2013qr ; Ikeda:2013vwa ; Bicudo:2015kna ; Bicudo:2015vta ; Peters:2015tra ; Peters:2016isf ; Francis:2016hui ; Bicudo:2017szl ; Chen:2017rhl ; Azizi:2018mte ; Bicudo:2019mny ; Wang:2019xzt ; Lu:2020qmp ; He:2020jna ; Cheng:2020nho ; Albuquerque:2020ugi ; Guo:2021mja ; Chen:2022asf for comprehensive reviews). These studies have deepened our understanding of the exotic hadron spectrum and highlighted the importance of investigating the interactions of dimeson systems, particularly those involving doubly heavy quarks.
@@
In this work, we investigate the interactions of $S$-wave $D\bar{B}$, $D\bar{B}^{*}$, and $D^{*}\bar{B}^{*}$ systems using chiral effective field theory (ChEFT) in the heavy hadron formalism. ChEFT, a low-energy effective theory of QCD, has proven to be a powerful tool for studying hadronic interactions, particularly in the context of nucleon-nucleon systems Machleidt:2011zz ; Hammer:2019poc . By incorporating heavy-quark and chiral symmetries, ChEFT provides a robust framework for analyzing the interactions of heavy mesons. Previous studies have applied ChEFT to systems such as $D^{(*)}D^{(*)}$ Xu:2017tsr ; Wang:2022jop , $\bar{B}^{(*)}\bar{B}^{(*)}$ Wang:2018atz ; Abreu:2022sra , and hidden-charm systems Xu:2021vsi , as well as to the interactions of heavy pentaquark molecular states Meng:2019ilv ; Wang:2019ato ; Wang:2022ztm . These investigations have demonstrated the effectiveness of ChEFT in describing the low-energy dynamics of heavy hadrons.
This paper is organized as follows: In Sec. II, we introduce the chiral effective Lagrangians with $SU(2)$ flavor symmetry used in our calculations. In Sec. III, we present the scattering amplitudes and derive the effective meson-meson potentials. In Sec. IV, we analyze the numerical results, examining the behaviors of the effective potentials and searching for possible bound states by solving the Schrödinger equation. We also examine the coupled-channel effects. We discuss the scattering $T$-matrix and extract physical quantities such as the scattering rate, scattering length, and effective range. Finally, in Sec. V, we summarize our findings and discuss their implications for future studies of doubly heavy tetraquark systems.
@@ II.2 Lagrangians at next-to-leading order
The low-energy constants appearing in the above equations are split into finite and infinite parts. The infinite parts are used to cancel the divergences of the loop contributions at NLO. The finite parts also contribute but are not determined here due to the lack of data input. Therefore, we just discard them.
@@
where $\mathcal{F}(\mathbf{q})$ is a regulator function. To avoid ultraviolet (UV) divergence in the integral, we use a Gaussian regulator function $\mathcal{F}(\mathbf{q})=\exp(-\mathbf{q}^{2n}/\Lambda^{2n})$ and an adequately large $n$ is adopted so that the powers generated by the regulator are beyond the order at which our calculation is performed so that it does not affect the accuracy at the given order Xu:2017tsr ; Wang:2018atz ; Xu:2021vsi ; Entem:2003ft ; Entem:2017gor . Therefore, we chose $2n>\epsilon$ for the contact terms at order $\epsilon$. Here, we take $n=2$. Based on the potentials in coordinate space, we can solve the Schrödinger equations to search for the possible bound states.
@@ III.1 $D\bar{B}$ systems
In the above amplitudes (19)-(29), the loop functions $J^{a/b}_{ij}(m,\omega)$, $J^{g/h}_{ij}(m,\omega_{1},\omega_{2})$, $J^{F}_{ij}(m_{1},m_{2},q)$, $J^{T/S}_{ij}(m_{1},m_{2},\omega,q)$, and $J^{B/R}_{ij}(m_{1},m_{2},\omega_{1},\omega_{2},q)$ are abbreviated as $J^{a/b}_{ij}$, $J^{g/h}_{ij}$, $J^{F}_{ij}$, $J^{T/S}_{ij}$, and $J^{B/R}_{ij}$, respectively. The variables $m,m_{1},m_{2}$ denote the pion mass, while the mass-splitting-dependent variables $\omega_{1}$ and $\omega_{2}$ for every amplitudes are collected in Tables 1-2. Here we define $\delta_{1}=M_{D^{*}}-M_{D}$ and $\delta_{2}=M_{\bar{B}^{*}}-M_{\bar{B}}$ as the $D^{*}$-$D$ and $\bar{B}^{*}$-$\bar{B}$ mass differences, respectively. We list the definitions of the loop functions in Appendix B, and the calculation procedures of these loop functions follow the Refs. Xu:2017tsr ; Xu:2021vsi .
Figure 1: LO contact, NLO contact and NLO TPE diagrams of the process $D\bar{B}\to D\bar{B}$. The thin solid, double thin solid, solid, thick solid, and dashed lines stand for the $\bar{B}$, $\bar{B}^{*}$, $D$, $D^{*}$ and a pion, respectively.
@@ III.2 $D\bar{B}^{*}$ systems
$\mathcal{M}_{a2.2}^{(2)}=$ $~{}4\frac{g_{2}^{2}}{f^{2}}(d-3)(d-2)AJ_{22}^{g}(\epsilon_{2}\cdot\epsilon^{*}_{4}),$ (33)
$\mathcal{M}_{a2.3}^{(2)}=$ $~{}4\frac{g_{1}^{2}}{f^{2}}(d-1)AJ_{22}^{g}(\epsilon_{2}\cdot\epsilon^{*}_{4}),$ (34)
@@
$\mathcal{M}_{a2.6}^{(2)}=$ $-4\frac{g_{1}g_{2}}{f^{2}}(d-3)(d-2)AJ_{22}^{h}(\epsilon_{2}\cdot\epsilon^{*}_{4}),$ (37)
$\mathcal{M}_{a2.7}^{(2)}=$ $-4\frac{g_{1}g_{2}}{f^{2}}(d-3)(d-2)AJ_{22}^{h}(\epsilon_{2}\cdot\epsilon^{*}_{4}),$ (38)
@@
$\mathcal{M}_{a2.10}^{(2)}=$ $~{}\frac{3g_{1}^{2}}{2f^{2}}(d-2)A\partial_{\omega}J_{22}^{b}(\epsilon_{2}\cdot\epsilon^{*}_{4}).$ (40)
@@
$\mathcal{M}_{c2.2}^{(2)}=$ $~{}i4\frac{g_{2}^{2}}{f^{4}}\{[A_{1}(q_{0}J_{22}^{S}+J_{24}^{S}+q_{0}J_{32}^{S}+J_{33}^{S})-A_{5}(q_{0}J_{11}^{S}+2q_{0}J_{22}^{S}+J_{24}^{S}+q_{0}J_{32}^{S}+J_{33}^{S})](q\cdot\epsilon_{2})(q\cdot\epsilon^{*}_{4})$
@@
$\mathcal{M}_{c2.5}^{(2)}=$ $~{}4\frac{g_{1}^{2}g_{2}^{2}}{f^{4}}A_{1}\{[J_{21}^{B}-\vec{q}^{2}J_{22}^{B}+(d+3)(J_{31}^{B}+J_{42}^{B})-2\vec{q}^{2}J_{32}^{B}-\vec{q}^{2}J_{43}^{B}](q\cdot\epsilon_{2})(q\cdot\epsilon^{*}_{4})-[\vec{q}^{2}(J_{31}^{B}$
$+J_{42}^{B})-(1+d)J_{41}^{B}](\epsilon_{2}\cdot\epsilon^{*}_{4})\},$ (45)
$\mathcal{M}_{c2.6}^{(2)}=$ $~{}4\frac{g_{1}^{2}g_{2}^{2}}{f^{4}}(d-3)A_{1}\{[J_{21}^{B}-\vec{q}^{2}J_{22}^{B}+(d+3)J_{31}^{B}-2\vec{q}^{2}J_{32}^{B}+(d+3)J_{42}^{B}-\vec{q}^{2}J_{43}^{B}](q\cdot\epsilon_{2})(q\cdot\epsilon^{*}_{4})$
$+[-\vec{q}^{2}J_{21}^{B}+\vec{q}^{4}J_{22}^{B}-(2d+1)\vec{q}^{2}(J_{31}^{B}+J_{42}^{B})+2\vec{q}^{4}J_{32}^{B}+(d+1)(d-2)J_{41}^{B}+\vec{q}^{4}J_{43}^{B}](\epsilon_{2}\cdot\epsilon^{*}_{4})\},$ (46)
$\mathcal{M}_{c2.7}^{(2)}=$ $~{}4\frac{g_{1}^{2}g_{2}^{2}}{f^{4}}A_{1}\{[J_{21}^{R}-\vec{q}^{2}J_{22}^{R}+(d+3)(J_{31}^{R}+J_{42}^{R})-2\vec{q}^{2}J_{32}^{R}-\vec{q}^{2}J_{43}^{R}](q\cdot\epsilon_{2})(q\cdot\epsilon^{*}_{4})-[\vec{q}^{2}J_{31}^{R}$
$-(1+d)J_{41}^{R}+\vec{q}^{2}J_{42}^{R}](\epsilon_{2}\cdot\epsilon^{*}_{4})\},$ (47)
$\mathcal{M}_{c2.8}^{(2)}=$ $~{}4\frac{g_{1}^{2}g_{2}^{2}}{f^{4}}(d-3)A_{1}\{[J_{21}^{R}-\vec{q}^{2}J_{22}^{R}+(d+3)J_{31}^{R}-2\vec{q}^{2}J_{32}^{B}+3J_{42}^{B}-\vec{q}^{2}J_{43}^{B}](q\cdot\epsilon_{2})(q\cdot\epsilon^{*}_{4})+[-\vec{q}^{2}J_{21}^{R}$
$+\vec{q}^{4}J_{22}^{R}-(2d+1)\vec{q}^{2}(J_{31}^{R}+J_{42}^{R})+2\vec{q}^{4}J_{32}^{R}+(d+1)(d-2)J_{41}^{R}+\vec{q}^{4}J_{43}^{R}](\epsilon_{2}\cdot\epsilon^{*}_{4})\},$ (48)
@@
. $I=1$ $I=0$ $A$ $A$ $\omega_{1}$ $\omega_{2}$ $A_{a2.1}$ $\frac{1}{4}(3D_{a}-E_{a})$ $\frac{3}{4}(D_{a}+E_{a})$ $\delta_{2}$ $\delta_{2}$ $A_{a2.2}$ $\frac{1}{4}(3D_{a}-E_{a})$ $\frac{3}{4}(D_{a}+E_{a})$ 00 00 $A_{a2.3}$ 00 00 $-\delta_{1}$ $-\delta_{1}$ $A_{a2.4}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{-3}{4}(D_{b}-3E_{b})$ 00 $-\delta_{1}$ $A_{a2.5}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{-3}{4}(D_{b}-3E_{b})$ $-\delta_{1}$ 00 $A_{a2.6}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{-3}{4}(D_{b}-3E_{b})$ $\delta_{2}$ $-\delta_{1}$ $A_{a2.7}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{-3}{4}(D_{b}-3E_{b})$ $-\delta_{1}$ $\delta_{2}$ $A_{a2.(8+9)}$ $D_{a}+E_{a}$ $D_{a}-3E_{a}$ 00 $\delta_{2}$ $A_{a2.10}$ $D_{a}+E_{a}$ $D_{a}-3E_{a}$ $-\delta_{1}$ 00
Figure 2: LO contact, NLO contact and NLO TPE diagrams of the process $D\bar{B}^{*}\to D\bar{B}^{*}$. The thin solid, double thin solid, solid, thick solid, and dashed lines stand for the $\bar{B}$, $\bar{B}^{*}$, $D$, $D^{*}$, and a pion, respectively.
@@
       $A_{c2.2}$        $\frac{i}{8}$        $\frac{-i}{8}$        00        $-\frac{3i}{8}$        $\frac{3i}{8}$        00        00        00
       $A_{c2.3}$        $\frac{i}{8}$        $-\frac{i}{8}$        00        $-\frac{3i}{8}$        $\frac{3i}{8}$        00        $\delta_{2}$        00
@@
       $A_{c2.5}$        $\frac{1}{16}$        00        00        $\frac{9}{16}$        00        00        $-\delta_{1}$        00
       $A_{c2.6}$        $\frac{1}{16}$        00        00        $\frac{9}{16}$        00        00        $-\delta_{1}$        $\delta_{2}$
       $A_{c2.7}$        $\frac{5}{16}$        00        00        $-\frac{3}{16}$        00        00        $-\delta_{1}$        00
       $A_{c2.8}$        $\frac{5}{16}$        00        00        $\frac{-3}{16}$        00        00        $-\delta_{1}$        $\delta_{2}$